Rigo:

A rectangular room is 4 times as long as it is wide, and its perimeter is 50 meters. Find the dimension of the room.

Rigo:

idk guys

jhonyy9:

 P = 50 P - perimeter how you calcule the perimeter of a rectangle ?

Rigo:

idk

jhonyy9:

what mean perimeter ?

Rigo:

find the demension of the room

Rigo:

like w=x meters or L= x meters

jhonyy9:

yes you can calcule this x and y values than you know what mean perimeter of a rectangle

Rigo:

idk how tho

jhonyy9:

perimeter of a rectangle = 2 length + 2 width

Rigo:

ok thx

jhonyy9:

pleas make the calcules and i like check it

Rigo:

w=8 L=32

jhonyy9:

and using these value you get for perimeter 50 ?

Rigo:

yah

jhonyy9:

2*8=16 2*32=64 16+64 = ?

jhonyy9:

no

jhonyy9:

4x +4x +x+x = 50 x = ?

Rigo:

my bad, ik the peremiter, i need the width and lenght

jhonyy9:

so from these above wrote equation you need calcule the x value and so will can calcule easy the length and width

Rigo:

oh